1.   A certain type of elementary particle travels at a speed of 2.70 𝑥 108 𝑚/𝑠. At this speed, the average lifetime is measured to be 4.76 𝑥 10−6 𝑠. What is the particle’s lifetime at rest?



1
Expert's answer
2022-04-04T09:06:48-0400

Speed of particle is "v=2.7\times 10^8 m/s"

Measured lifetime of particle is "\Delta t=4.76\times 10^{-6} s"

The average lifetime of particle at rest is

"\Delta t_0=\Delta t \sqrt{1- (\frac{v}{c})^2}"

"=(4.76\times 10^{-6} s)\sqrt{1- (\frac{2.7 \times10^8}{3.0 \times10^8})^2}"

"=2.075 \times 10^{-6} s"